Here's the other part!

(ii) Find the molar concentration of the sulfuric acid from the titration result:
Complete the calculation given below:
No. of moles of KOH in 25.00 cm3 =
= ………………………… (1 mark)
 no. of moles H2SO4 in 15.00 cm3 =
= ………………………… (1 mark)
 molar concentration of H2SO4 =
(Include units in your answer) = ………………………… (2 marks)

(Total 10 marks)
1
Expert's answer
2021-07-17T04:44:54-0400

Relative Molecular Mass (RMM) of

KOH is 56.11 g/mole


No. of moles of KOH in 250 cm3"= \frac{0.95}{56.11}=0.0169moles"


2KOH (aq) + H2SO4 (aq) → K2SO4 (aq) + 2H2O (l)


Moles of "H_2SO_4" "=\frac{0.0169}{2}=0.00845moles"


Molarity ="\frac{0.00845}{0.015}=0.563M"